Sam is 4 years older than her brother Jaems and one-half the age of her cousain Robert. If the sum of their ages is 60, how old is James
Answer by stanbon(75887)

Sam is 4 years older than her brother Jaems and one-half the age of her cousain Robert. If the sum of their ages is 60, how old is James
-------------------
Equations:
s = j + 4
-----
s = (1/2)r
----
s + j + r = 60
----
Substitute for "j" and "r" to get:
s + s-4 + 2s = 60
4s = 64
s = 16
----
James = s-4 = 16-4 = 12 yrs old
===================================
cheers,
Stan H.